# Splittable Spanning Trees and Balanced Forests in Dense Random Graphs

### v2

@@ 3. Challenges when allowing slack
  1. (1)
    is a Hamiltonian induced subgraph of the triangular lattice
  2. (2)
    contains exactly $n$ partitions into three pieces of size $n$, namely the partitions of the $3n$ Hamiltonian paths
  3. (3)
    is such that $\tau^{0}(G_{n})/\tau^{1}(G_{n})$ is exponentially small.
Corollary 3.1.
@@
Figure 1. Graph described in Theorem 3 with $n=14$: (a) An almost balanced partition, with piece sizes $\{13,14,15\}$ and (b) an example of a balanced partition, with piece sizes $\{14,14,14\}$.
@@
We can now proceed with the construction of the family of graphs discussed in Theorem 3.
Proof of Theorem 3.
@@
Now consider any balanced partition $P$ of $G_{n}$ into induced subgraphs $P_{1},P_{2},P_{3}$. Consider the subgraph $L_{i}$ of $P_{i}$ induced by the maximal ladders in $P_{i}$. By the discussion above, $P_{i}$ consists of components of $L_{i}$, connected in a chain by paths. Therefore, $\tau(P_{i})$ = $\tau(L_{i})$. By Lemma 3.1 it will be enough to show that the monochromatic ladders of $P$ contain $3n-\Omega(n)$ vertices total.
There are four ladders and only three pieces, so either there are fewer than four monochromatic ladders, in which case we are done, or one piece $P_{i}$ of the partition contains two ladders. Those two ladders must be adjacent in the clockwise traversal of the Hamiltonian cycle. Any pair of adjacent ladders in $G_{n}$ contain $n+\Omega(n)$ total vertices, of which $P_{i}$ must be missing $\Omega(n)$. ∎
Proof Of Corollary.
This implies that when using ReCom (or its variants) without slack, no move is possible, as any piece’s position along the Hamiltonian cycle determines the positions of the other two pieces. Hence, any move splitting and recombining two of them cannot modify the partition, so ReCom fails to make any nontrivial moves. As the graph can only be partitioned along the Hamiltonian cycle, there are only $n$ balanced partitions of the graph up to relabeling. We also see that when allowing any slack, we must allow the partitions with the smallest amount of slack: those with piece sizes n1,n,n+1n-1,n,n+1. Then any sample made from ReCom with slack is exponentially more likely to unbalanced than balanced, the example high weight unbalanced partition described in theorem 3 is clearly reachable as a ReCom move from a balanced partition along the Hamiltonian cycle. ∎
4. Splittability of Dense Graphs
Figure 2. Total spanning tree weights of all Partitions of size i,Nii,N-i on the Complete Graph with $N=8$ (a) and $N=16$ (b).
@@ 4. Splittability of Dense Graphs
As Section 3 demonstrates, even the smallest amount of slack can cause balanced partitions to be exponentially suppressed at stationarity. This potentially poses a problem for any algorithms sampling spanning trees until a splittable one is found. In this section we focus on Algorithms 1 and 2, that are efficient only when balanced partitions (or balanced forests) are not too unlikely compared to the weight when they balance is not required.
4.1. Splittability of the complete graph
We now show that in the complete graph a polynomial fraction of spanning trees are splittable. This is equivalent to showing that a polynomial fraction of partitions of the complete graph are balanced (Lemma 2.2), which implies that we can find a balanced partition of the complete graph in polynomial time by generating random spanning trees and using rejection sampling.
@@ 4.1. Splittability of the complete graph
In the next subsection we will extend this result to dense random graphs. We first prove our result on the complete graph because we will use it in the proof of the stronger (more general) result that follows. Recall Theorem 1, given as stated in the introduction:
Theorem 1.
@@
Proof.
By Cayley’s Theorem[Cayley] the number of spanning trees in $K_{N}$ is $N^{N-2}$. Each splittable spanning tree can be split uniquely into balanced pieces (Lemma 2.1); therefore, each one is uniquely defined by the partition it creates, the spanning trees on the $K_{n}$s in those partitions, and the k1k-1 edges between partitions. The number of unlabeled $k$-partitions of $N$ vertices is
@@
Dividing by the number of spanning trees in $K_{N}$ and applying Stirling’s formula  [abramowitz] to each of the factorials in this expression, the probability a spanning tree is splittable is
$\frac{\tau_{k}(K_{n})}{{(kn)}^{kn-2}}\$$=\ \left((1+O(\frac{1}{n}))\sqrt{2\pi n}(\frac{e}{n})^{n})^{k}~(1+O(\frac{1}{k})\right)$
@@
Setting $C_{k}=(1+o(k))(\frac{e}{\sqrt{2\pi}})^{k}$, the theorem follows. ∎
@@
4.2. Dense random graphs
We now show that in a dense random graph a polynomial fraction of spanning trees are splittable. This is equivalent to showing that a polynomial fraction of partitions of a dense random graph are balanced (Lemma 2.2), which implies that we can find a balanced partition of a dense random graph in polynomial time by generating random spanning trees and using rejection sampling.
While each spanning tree (with the nodes labeled) is equally likely to appear in a random graph, it is not true that any two spanning trees are equally likely to be selected as a random spanning tree of a random graph. The probability that any given tree is selected as a random spanning tree is proportional to $\mathbf{E}[1/\tau(G)\,|\,T\in G]$, which varies for different trees. Trees that occur in graphs that contain many other trees are less likely to be selected.
@@ 4.2. Dense random graphs
Recall Theorem 2, which states:
Theorem 2.
@@
For random connected $G$, the probability that a random spanning tree is connected is the expected value of this ratio,
@@
from Theorem 1.
Grimmett [grimmett_upper_1976] has shown that $\tau_{max}(m,N)\leq\frac{1}{N}(\frac{2m}{N-1})^{N-1}=p^{N-1}N^{N-2}$; therefore, the probability that our tree $T_{0}$ is $k$-splittable is at least
@@
which proves the theorem. ∎
@@
Proof.
Because the probability of being connected increases with $p$ we can prove the lemma for an arbitrary random graph $G$ in the $G(N,p)$ model. We will show that for each $j\leq p\binom{N}{2}$, $G$ is at least half as likely to have $p\binom{N}{2}+j$ edges as $p\binom{N}{2}-j$ edges. Letting $M\ =\ \binom{N}{2}$ and $m=pM$,
@@
For any fixed $m\geq p^{\prime}\binom{N}{2}$, if we condition on $G$ having $m$ edges, then both $G(N,m)$ and $G(N,p)$ have the uniform distribution over connected graphs. We now apply Theorem 4 for each such $m$. Recalling that the bound in Theorem 4 is an increasing function of $m/\binom{N}{2}$, the probability that $T$ is $k$-splittable is at least
@@
where we have absorbed the constants $1/3$ and $p^{\prime}/p$ into $C_{k}$. The theorem follows. ∎
Note that Theorem 2] follows immediately from Theorems 4 and 5.
5. Conclusions and Open Problems
@@
We also observe a link between random spanning trees of random graphs and minimum spanning trees with random weights in the complete graph. We conjecture that random spanning trees of random graphs will have smaller diameter than random spanning trees of the complete graph, as this occurs on random minimum spanning trees [addarioberry2017] and in our empirical results on random graphs.

### v3

@@ 3 Obstacles, even with slack
  1. 1.
    is a Hamiltonian induced subgraph of the triangular lattice,
  2. 2.
    contains exactly $n$ partitions into three pieces of size $n$, namely the partitions of the $3n$ Hamiltonian paths,
  3. 3.
    is such that $\tau^{0}(G_{n})/\tau^{1}(G_{n})$ is bounded above by an exponential $e^{-\Theta(n)}$.
Corollary 1
@@
Figure 1: Graph described in Theorem 3.1 with $n=14$: (a) An almost balanced partition, with piece sizes $\{13,14,15\}$ and (b) an example of a balanced partition, with piece sizes $\{14,14,14\}$.
@@
We can now proceed with the construction of the family of graphs discussed in Theorem 3.1.
Proof(Proof of Theorem 3.1)
@@
Now consider any balanced partition $P$ of $G_{n}$ into induced subgraphs $P_{1},P_{2},P_{3}$. Consider the subgraph $L_{i}$ of $P_{i}$ induced by the maximal ladders in $P_{i}$. By the discussion above, $P_{i}$ consists of components of $L_{i}$ connected in a chain by paths. Therefore, $\tau(P_{i})$ = $\tau(L_{i})$. By Lemma 5 it will be enough to show that the monochromatic ladders of $P$ contain $3n-\Omega(n)$ vertices total.
There are four ladders and only three pieces, so either there are fewer than four monochromatic ladders, in which case we are done, or one piece $P_{i}$ of the partition contains two ladders. Those two ladders must be adjacent in the clockwise traversal of the Hamiltonian cycle. Any pair of adjacent ladders in $G_{n}$ contain $n+\Omega(n)$ total vertices, of which $P_{i}$ must be missing $\Omega(n)$.
Proof(Proof of Corollary 1)
When using ReCom (or its variants) without slack on the family of graphs constructed for the proof of Theorem 3.1, Property (2) implies that no move is possible, as any piece’s position along the Hamiltonian cycle determines the positions of the other two pieces. Hence, any move splitting and recombining two of them cannot modify the partition, and ReCom fails to make any nontrivial moves. As the graph can only be partitioned along the Hamiltonian cycle, there are only $n$ balanced partitions of the graph up to relabeling. We also see that when allowing any slack, we must allow the partitions with the smallest amount of slack: those with piece sizes n1,n,n+1n-1,n,n+1. Then any sample made from ReCom with slack is exponentially more likely to be unbalanced than balanced, as the example high-weight unbalanced partition described in theorem 3.1 is clearly reachable as a ReCom move from a balanced partition along the Hamiltonian cycle.
4 Splittability of Dense Graphs
Figure 2: Total spanning tree weights of all Partitions of size i,Nii,N-i on the Complete Graph with $N=8$ (a) and $N=16$ (b).
@@ 4 Splittability of Dense Graphs
As Section 3 demonstrates, even the smallest amount of slack can cause balanced partitions to be exponentially suppressed at stationarity. This potentially poses a problem for any algorithms sampling spanning trees until a splittable one is found. In this section we focus on Algorithms 1 and 2, that are efficient only when balanced partitions (or balanced forests) are not too unlikely compared to the total weight where balance is not required.
4.1 Splittability of the complete graph
We now show that in the complete graph a polynomial fraction of spanning trees are splittable. This is equivalent to showing that a polynomial fraction of partitions of the complete graph are balanced (Lemma 2), which implies that we can find a balanced partition of the complete graph in polynomial time by generating random spanning trees and using rejection sampling.
@@ 4.1 Splittability of the complete graph
In the next subsection we will extend this result to dense random graphs. We first prove our result on the complete graph because we will use it in the proof of the stronger (more general) result that follows. Recall Theorem 4.1, given as stated in the introduction:
Theorem 4.1
@@
Proof
By Cayley’s Theorem[11] the number of spanning trees in $K_{N}$ is $N^{N-2}$. Each splittable spanning tree can be split uniquely into balanced pieces (Lemma 1); therefore, each one is uniquely defined by the partition it creates, the spanning trees on the $K_{n}$s in those partitions, and the k1k-1 edges between partitions. The number of unlabeled $k$-partitions of $N$ vertices is
@@
Dividing by the number of spanning trees in $K_{N}$ and applying Stirling’s formula  [1] to each of the factorials in this expression, the probability a spanning tree is splittable is
$\frac{\tau_{k}(K_{kn})}{{(kn)}^{kn-2}}\$$=\ \left((1+O(\frac{1}{n}))\sqrt{2\pi n}(\frac{e}{n})^{n})^{k}~(1+O(\frac{1}{k})\right)$
@@
Setting $C_{k}=(1+o(k))(\frac{e}{\sqrt{2\pi}})^{k}$, the theorem follows.
@@
4.2 Dense random graphs
We now show that in a dense random graph a polynomial fraction of spanning trees are splittable. This is equivalent to showing that a polynomial fraction of partitions of a dense random graph are balanced (Lemma 2), which implies that we can find a balanced partition of a dense random graph in polynomial time by generating random spanning trees and using rejection sampling.
While each spanning tree (with the nodes labeled) is equally likely to appear in a random graph, it is not true that any two spanning trees are equally likely to be selected as a random spanning tree of a random graph. The probability that any given spanning tree is selected as a random spanning tree is proportional to $\mathbf{E}[1/\tau(G)\,|\,T\in G]$, which varies for different spanning trees. Spanning trees that occur in graphs that contain many other spanning trees are less likely to be selected.
@@ 4.2 Dense random graphs
Our main theorem is as follows:
Theorem 4.2
@@
For random connected $G$, the probability that a random spanning tree is $k$-splittable is the expected value of this ratio,
@@
from Theorem 4.1.
Grimmett [15] has shown that $\tau_{max}(m,N)\leq\frac{1}{N}(\frac{2m}{N-1})^{N-1}=p^{N-1}N^{N-2}$; therefore, the probability that our tree $T_{0}$ is $k$-splittable is at least
@@
which proves the theorem.
@@
Proof
Let $C$ be the event that $G\in G(N,p)$ is connected and $L$ be the event that $G$ has at least $p\binom{N}{2}$ edges. We want to prove that $P(L|C)=P(L)P(C|L)/P(C)\geq 1/3$. Because $P(C|L)/P(C)\geq 1$, we can prove that $P(L)\geq 1/3$ for an arbitrary random graph $G$ in the $G(N,p)$ model. We will show that for each $j\leq p\binom{N}{2}$, $G$ is at least half as likely to have $p\binom{N}{2}+j$ edges as $p\binom{N}{2}-j$ edges. Letting $M\ =\ \binom{N}{2}$ and $m=pM$,
@@
For any fixed $m\geq p^{\prime}\binom{N}{2}$, if we condition on $G$ having $m$ edges, then both $G(N,m)$ and $G(N,p)$ have the uniform distribution over connected graphs. We now apply Theorem 4.3 for each such $m$. Recalling that the bound in Theorem 4.3 is an increasing function of $m/\binom{N}{2}$, the probability that $T$ is $k$-splittable is at least
@@
where we have absorbed the constants $1/3$ and $p^{\prime}/p$ into $C_{k}$. The theorem follows.
Note that Theorem 4.2 follows immediately from Theorems 4.3 and $4.$
5 Random Spanning Tree Diameters
One might ask how these algorithms compare in practice, so we include a table of values showing our simulations of the average diameter of random spanning trees on random and complete graphs, and minimum spanning trees of complete graphs with random values. These empirical results helped motivate our questions surrounding the questions of random structures on random graphs.
Each entry in the table is the mean diameter of random spanning trees, taken over $3000$ trials. In the minimum spanning tree (MST) case, a complete graph on $n$ vertices is given random weights uniformly in $[0,1]$, and the diameter of the minimum spanning tree is taken, and in the other cases, a $G(n,p)$ random graph is created each trial, and the diameter of a uniform random spanning tree is measured. Trials where the random graph was not connected are not included in the average.
We see that the smallest diameters occur on minimum spanning trees, and random spanning trees on $G(n,p)$ graphs with smaller values of $p$ have clearly smaller diameters as $p$ decreases. The excepted diameter of a random minimum spanning tree scales with $O(n^{1/3})$[2] while the expected diameter of a random spanning tree of $K_{n}$ scales with $O(n^{1/2})$[18], and the scaling limit is unknown for a random spanning tree of a random graph. We note that spanning trees in $G(n,1/2)$ graphs have close enough average diameter to $K_{n}$ that a difference is not observed in this data.
Table 1: Empirical average spanning tree diameters on different randomization models
[…]
6 Conclusions and Open Problems
@@
We also observe a link between random spanning trees of random graphs and minimum spanning trees with random weights in the complete graph. We conjecture that random spanning trees of random graphs will have smaller diameter than random spanning trees of the complete graph, as this occurs on random minimum spanning trees [2][18] and in our empirical results on random graphs.